# Graded lithium-ion battery pouch cells to homogenise current distributions and mitigate lithium plating

### v1

@@ Conclusions
A model for large format lithium-ion battery pouch cells with spatial grading across the plane was developed and validated against experimental data. An analytic solution for the optimal electrical resistance distribution across the plane of large format pouch cells to achieve a flat current distribution was derived. The main benefit of graded electrodes identified was to increase the allowable charging C-rate from which lithium plating could be avoided; the graded electrodes increased the safe charging rate from 2.75C to 5.5C. By contrast, only a 1.2% increase in charging capacity was achieved for 4C charging. These results indicate the potential of in-the-plane graded electrode microstructures for enabling fast charging of large format pouch cells that by limiting the spatial heterogeneities.
@@ Acknowledgements
$U(\text{SoC},T)=\frac{dU}{dT}\ \left(T-T_{\scriptscriptstyle\text{ref}}\right)+0.0027\ T\ I_{sign}+3.382+0.0047\ \text{SoC}+1.627\ \exp(-81.163\ \text{SoC}^{1.0138})\\ +7.6445\times 10^{-8}\ \exp(25.36\ \text{SoC}^{2.469})-8.441\times 10^{-8}\ \exp(25.262\ \text{SoC}^{2.478})-3.3852+3.2585,$ (A1)
@@
$V(y,z,0)$ $=V_{0}=U_{0}+\ R_{0}\ I_{\scriptscriptstyle\text{app}}/A_{\text{tab}},\ \text{on}\ \Omega\ \text{and}\ \partial\Omega,$ (A2c)
@@
$i_{x}$ Current density through electrode thickness A///m2
$I$ Current density in the current collector A///m2
@@
$R$ Cell series resistance area $\Omega$m2
@@
$C_{\scriptscriptstyle\text{cell}}$ Capacity of cell (20 Ah) per cell area A s///m2 20 ×$\times 3600/A_{\scriptscriptstyle\text{cell}}$ Experiment
$C_{1}$ Capacitance of RC-pair F 3.49×$\times 10^{4}$ Fitting
$C_{2}$ Capacitance of RC-pair F 1.11×$\times 10^{4}$ Fitting
[…]
@@
$L_{\scriptscriptstyle\text{x}}$ Electrode thickness including $n=42$ layers (110 $\mu$m without $n$) m 0.0046 Experiment
@@
$R_{0}$ Series resistance of Uniform cell $\Omega$ m2 1.5×$\times 10^{-3}\times A_{\scriptscriptstyle\text{cell}}$ Fitting
$R_{1}$ Resistance of RC-pair $\Omega$ m2 9.03×$\times 10^{-4}\times A_{\scriptscriptstyle\text{cell}}$ Fitting
$R_{2}$ Resistance of RC-pair $\Omega$ m2 1.80×$\times 10^{-4}\times A_{\scriptscriptstyle\text{cell}}$ Fitting
@@
$\lambda_{\scriptscriptstyle\text{eff}}$ Effective thermal conductivity W///m///K 4.5 Fitting
$\rho$ Heat capacity J///K///m3 125530 Fitting
$\sigma_{\scriptscriptstyle\text{cp}}$ Electronic conductivity of copper positive current collector S/m $5.96\times 10^{7}$ Experiment
$\sigma_{\scriptscriptstyle\text{cn}}$ Electronic conductivity of aluminium negative current collector S///m $3.77\times 10^{7}$ Experiment
Figure A1: Comparison between the graded and uniform models during cycling at 4C and an applied current of 80 A. The current and temperature responses with time are shown, including the minimum, average, and maximum temperatures across the plane of the pouch cell. The voltage is similar between the two models.

### v2

@@ Conclusions
A model for large format lithium-ion battery pouch cells with spatial grading across the plane was developed and validated against experimental data. An analytical solution for the optimal electrical resistance distribution across the plane of large format pouch cells to achieve a flat current distribution was derived. The main benefit of graded electrodes identified was to increase the allowable charging C-rate from which lithium plating could be avoided; the graded electrodes increased the safe charging rate from 2.4C to 4.3C. By contrast, only a 1.2% increase in charging capacity was achieved for 4C charging. These results indicate the potential of in-the-plane graded electrode microstructures for enabling fast charging of large format pouch cells that by limiting the spatial heterogeneities. Future work will focus on experiments to fabricate graded electrodes for LFP pouch cells and validate the results of this modelling work.
@@ Acknowledgements
$U(\text{SoC},T)=\frac{dU}{dT}\ \left(T-T_{\scriptscriptstyle\text{ref}}\right)+0.0027\ T\ I_{\text{sign}}+3.382+0.0047\ \text{SoC}+1.627\ \exp(-81.163\ \text{SoC}^{1.0138})\\ +7.6445\times 10^{-8}\ \exp(25.36\ \text{SoC}^{2.469})-8.441\times 10^{-8}\ \exp(25.262\ \text{SoC}^{2.478})-3.3852+3.2585,$ (A1)
@@
$V(y,z,0)$ $=V_{0}=U_{0}+\ R_{0}\ I_{\scriptscriptstyle\text{app}},\ \text{on}\ \Omega\ \text{and}\ \partial\Omega,$ (A2c)
@@
$i_{x}$ Current density through electrode thickness A m-2
$I$ Current density in the current collector A m-2
@@
$I_{\text{sign}}$ Sign of the current
$R$ Cell series resistance $\Omega$
@@
$n_{\text{layers}}$ Number of electrode layers in cell 42 Experiment
$C_{\scriptscriptstyle\text{cell}}$ Capacity of cell (20 Ah) per cell area A s 20 ×$\times 3600$ Experiment
$C_{1}$ Capacitance of RC-pair F 2.79×$\times 10^{4}$ Fitting
$C_{2}$ Capacitance of RC-pair F 8.89×$\times 10^{3}$ Fitting
$h$ Heat transfer coefficient W m-2 K-1 13.2 Fitting
$h_{\scriptscriptstyle\text{tab}}$ Heat transfer coefficient of tabs W m-2 K-1 51.58 Fitting
$h_{\scriptscriptstyle\text{faces}}$ Effective heat transfer coefficient of current collector faces W m-2 K-1 210×$\times L_{x}$ Fitting
@@
$L_{\scriptscriptstyle\text{x}}$ Electrode thickness m 110 ×$\times 10^{-6}$ Experiment
@@
$R_{0}$ Series resistance of Uniform cell $\Omega$ 1.5×$\times 10^{-3}$ Fitting
$R_{1}$ Resistance of RC-pair $\Omega$ 1.10×$\times 10^{-3}$ Fitting
$R_{2}$ Resistance of RC-pair $\Omega$ 2.25×$\times 10^{-4}$ Fitting
@@
$\lambda_{\scriptscriptstyle\text{eff}}$ Effective thermal conductivity W m-1 K-1 4.5 Fitting
$\rho$ Volumetric heat capacity J K-1 m-3 1.35 ×$\times 10^{5}$ Fitting
$\sigma_{\scriptscriptstyle\text{cn}}$ Electronic conductivity of copper negative current collector S m-1 $5.96\times 10^{7}$ Experiment
$\sigma_{\scriptscriptstyle\text{cp}}$ Electronic conductivity of aluminium positive current collector S m-1 $3.77\times 10^{7}$
Figure A1: Comparison between the graded and uniform models during cycling at 4C and an applied current of 80 A. The current and temperature responses with time are shown, including the minimum, average, and maximum temperatures across the plane of the pouch cell.